# MMSR: Symbolic Regression is a Multimodal Task

### v4

@@ 3.2 Training Objective
Black-box 0.9913 0.9655 0.9028
Average 0.9932 0.9633 0.9328
4 Experiments
4.1 Dataset
In order to test the performance of MMSR and baselines, we conducted comparative analyses using prevalent datasets, designated as ’Nguyen’, ’Keijzer’, ’Korns’, ’Constant’, ’Livermore’, ’Vladislavleva’, ’R’,’ Jin’, ’Neat’,’ Others’, ’Feynman’, ’Strogatz’ and ’Black-box’. The three datasets Feynman ’, ’Strogatz’ and ’Black-box’ are from the SRbetch dataset. These datasets contain nearly 400 expressions in total. We believe that they can accurately and comprehensively test the performance of each algorithm.
4.2 Compared with State-of-the-arts
Figure 6: From the above five plots, we can clearly see that the performance of the algorithm increases as the size of the training data increases. However, under the same data size, the model using the one-step strategy is better than the model using the two-step strategy. And the larger the amount of data, the more obvious this trend. The results for the remaining datasets are in Appendix Figures 9,10
  • SymbolicGPT: The classical algorithm for pre-trained models, which treats the symbolic regression problem as a translation problem. Treat a ‘single letter’ (e.g. ‘s’, ‘i’, ‘n’) as a token. Generating expressions sequentially.
  • NeSymReS: Based on SymbolicGPT, the algorithm treated each ‘operator’ (e.g. ‘sin’, ‘cos’) as a token. The sequence of expressions is generated in turn. The setup is more reasonable.
  • End2End: End2End further encodes the constants so that the model generates the constants directly when generating the expression, abandoning the constant placeholder ‘C’ used by the previous two algorithms.
  • TPSR: One uses a pre-trained model as a policy network to guide the MCTS process, which greatly improves the search efficiency of MCTS.
The radar chart in Figure.1 shows the performance comparison between our algorithm and four baselines intuitively. From the figure, we can find that MMSR has a better fitting performance than the four baselines.
4.3 Ablation Study
To verify whether the introduction of contrastive learning plays a positive role in the algorithm performance, we conduct the ablation experiment as follows. On the premise of ensuring that the training data and other factors are the same, According to Equation 2, we did three sets of experiments, in which $\lambda_{1}$ and $\lambda_{2}$ were always equal to 1, and $\lambda_{3}$ was set to 0.0, 0.1, and 1.0 in turn. The specific experimental results are shown in Table 1.
This can be clearly seen from the table. Other conditions being equal, as the proportion of contrastive learning loss ($\mathcal{L_{CON}}$) decreases gradually, the performance of the algorithm also decreases gradually. In addition, the overall performance with the contrastive loss ($\lambda_{3}$ = 1.0 and 0.1) is better than without the contrastive loss ($\lambda_{3}$ = 0.0). The contrastive loss makes the matching data features and skeleton features closer together in the feature space, and more importantly, it makes the non-matching data features and skeleton features further apart. In this way, the latter two features can be better fused.
4.4 Effects of modality alignment
To verify whether the extracted features before and after the introduction of contrastive learning are as expected: Data and skeleton features that match each other are close to each other in the feature space, whereas those that do not match are far away from each other. We first randomly generate 50 expression samples, feed these samples into the model without contrastive learning, and then obtain 50 pairs of features for the data and skeleton. We calculate the cosine distance between the extracted 50 data features and 50 skeleton features and then obtain the 50*50 cosine distance matrix A. Finally, we normalize the matrix A and visualize it to get fig 3(a). Similarly, we feed these samples into a model that uses contrastive learning, resulting in heatmap 3(b).
From figure 4, we can see that there is indeed a large change in the extracted features before and after the introduction of contrastive learning. In Figure 3(a), we can see that the data on the diagonal is not significantly smaller than the other positions, indicating that the matching data features and skeleton features are not the closest to each other in the feature space. On the contrary, we can clearly find in Figure 3b that the data on the diagonal is significantly smaller than the other positions, which indicates that our contrastive learning plays a role. The features of the two modalities, data and skeleton, have been well aligned.
4.5 Impact of the size of the training data
To test the impact of pre-training data size on model test performance, we train our MMSR model on increasingly larger datasets. More specifically, we use datasets consisting of 10K, 100K, 1M, 5M, and 10M samples to train a model respectively. Except for the amount of data, every other aspect of training is the same as described above, and all are trained for the same epochs.
@@ 4.5 Impact of the size of the training data
From the trend in Figure 5, we can see that the performance of the model has been greatly improved as the data set continues to grow. We can see that when the data size reaches 5M, the average $R^{2}$ has reached more than 0.9, and when the data size reaches 10M, The model achieves an average $R^{2}$ of over 0.98 on each dataset. Although the growth rate of average $R^{2}$ is relatively slow with the increase of data, it is still increasing. Therefore, it can be inferred that as we continue to increase the size of the training data, the performance of the model will still reach higher levels.
4.6 Effect of the introduction method of contrastive learning
It is very important to use contrastive learning for modality alignment in multi-modality. At present, there are two methods to introduce contrastive learning:
@@ 4.6 Effect of the introduction method of contrastive learning
  • One step. The contrast loss is used as a part of the total loss, and the feature extraction module and the feature fusion module are trained at the same time.
@@
  • Tow step. First, train a feature extraction module with contrastive learning, then freeze it, and then train a feature fusion module separately.
@@
Figure.3 shows the comparison plot between the One-step strategy and the Tow-step strategy. To prove the superiority of our adopted strategy, we conduct comparative experiments on all datasets. From figure 6, we can find that in terms of contrastive learning introduction method, the One-step strategy we adopted is slightly better than the Tow-step strategy in terms of effect. As the size of the dataset increases, this advantage becomes more and more obvious.
We believe that the specific reason for this phenomenon is that in the One-step strategy, contrastive loss and other losses are trained together. In the process of training, the feature extraction module and the feature fusion module are interrelated. This requires the feature extraction module not only to make the extracted data and skeleton features close in the feature space but also to make the extracted features better adapt to the later feature fusion module. So that the overall loss is smaller. In short, the one-step strategy can make each module better wearing in each other in the process of training together.
@@ 5 Discussion and conclusion
In this paper, we propose a symbolic regression algorithm, MMSR. We solve the SR problem as a purely multimodal problem and obtain decent experimental results. This also proves that it is practical to treat the SR problem as a multimodal task. Specifically, we treat the data and expression skeleton as two different modalities (e.g. image and text). For the data part, we use SetTransformer for the feature extractor, and for the expression skeleton part, we use the first few layers of the transformer as the feature extractor. Specifically, we introduce contrastive learning for modal feature alignment for better feature fusion. The ablation experiments on more than a dozen datasets show that the introduction of contrastive learning has a great improvement in the performance of the algorithm. In addition, we adopt the One-step scheme(the contrastive loss is trained along with other losses during training) in MMSR. Our experiments show that the One-step scheme is more competitive than the Tow-step scheme, which first trains a feature extraction module with contrastive loss and then trains the feature fusion module separately. In MMSR, alleviating expressions may exhibit high token-level similarities but are suboptimal with respect to equation-specific objectives such as fitting accuracy. Our decoder not only generates the sequence of symbols that make up the skeleton of an expression but also directly generates a set of constants. This set of constants is the same length as the symbol sequence, we take the constant with the same index as the placeholder ’C’ as the value at ’C’, and then the constants are further refined with BFGS to obtain the final expression.
Our algorithm is expected to open new application prospects of multimodal technology. It promotes more multimodal correlation techniques to be applied to the field of symbolic regression. We can even use symbolic regression techniques to re-feed multi-modal domains.
@@
Next, we will continue to explore the deep fusion of multimodal techniques and symbolic regression techniques. It also attempts to solve some of the existing problems of MMSR mentioned above.
@@ Acknowledgements
This work was supported in part by the National Natural Science Foundation of China under Grant 92370117, in part by CAS Project for Young Scientists in Basic Research under Grant YSBR-090 and in part by the Key Research Program of the Chinese Academy of Sciences under Grant XDPB22
Appendix A Supplementary material on the effect of data size on performance
This section shows the performance change trend of MMSR on Strogatz, R, Korns, Constant, Livermore, Black-box, Neat and Others datasets with the increase of training dataset size. From the figure, we can also see that the performance of MMSR increases with the growth of data size. This means that MMSR can theoretically get better results if we have more datasets.
Figure 7: From the above five plots, we can see that as the size of the training data increases, the performance of the model on each dataset continues to improve. When the data size reaches 10M, the model has achieved satisfactory results on the selected data sets, and the average $R^{2}$ on each data set is greater than 0.98.
Figure 8: From the above five plots, we can see that as the size of the training data increases, the performance of the model on each dataset continues to improve. When the data size reaches 10M, the model has achieved satisfactory results on the selected data sets, and the average $R^{2}$ on each data set is greater than 0.98.
Appendix B Supplementary material on one-step and two-step strategies
This section shows the performance trends of the one-step strategy and the two-step strategy on Strogatz, R, Korns, Constant, Livermore, Black-box, Neat and Others with the increase of the size of the training dataset. It can also be seen from the figure that with the growth of the data scale, no matter which strategy is used, the performance of the algorithm increases. However, the one-step strategy consistently outperforms the two-step strategy, and this trend becomes more and more obvious as the size of the data increases.
Figure 9: From the above five plots, we can clearly see that the performance of the algorithm increases as the size of the training data increases. However, under the same data size, the model using the one-step strategy is better than the model using the two-step strategy. And the larger the amount of data, the more obvious this trend.
Figure 10: From the above three plots, we can clearly see that the performance of the algorithm increases as the size of the training data increases. However, under the same data size, the model using the one-step strategy is better than the model using the two-step strategy. And the larger the amount of data, the more obvious this trend.
@@ Appendix C Appendix: Test data in detail
Table 2,3,4 shows in detail the expression forms of the data set used in the experiment, as well as the sampling range and sampling number. Some specific presentation rules are described below
@@
Table 2: Specific formula form and value range of the three data sets Nguyen, Korns, and Jin.
@@
${}^{\prime}$ $4x_{1}^{4}+3x_{1}^{3}+2x_{1}^{2}+x$ U$(-1,1,20)$
${}^{\prime}$ $\sin(x_{1}^{2})\cos(x)-2$ U$(-1,1,20)$
${}^{\prime}$ $3]{x}$ U$(0,4,20)$
${}^{\prime\prime}$ $3]{x_{1}^{2}}$ U$(0,4,20)$
@@
Table 3: Specific formula form and value range of the three data sets neat, Keijzer, and Livermore.
@@
Table 4: Specific formula form and value range of the three data sets Vladislavleva and others.
@@
The empirical results from our investigation unequivocally affirm that MMSR possesses an exceptional ability to discern the underlying mathematical expressions from a constrained sample size. Notably, the $R^{2}$ values achieved were above 0.99 for a predominant portion of the equations, underscoring the algorithm’s remarkable accuracy in fitting these expressions. These findings decisively position MMSR as a potent tool for addressing complex problems within the domains of physics and mathematics. The broader implications of our study suggest that MMSR holds considerable promise for a wide range of applications across different fields. Detailed experimental results are presented in Table 5 and Table 6.
@@
I.6.20a $f=e^{-\theta^{2}/2}/\sqrt{2\pi}$ 0.9999
I.6.20 $f=e^{-\frac{\theta^{2}}{2\sigma^{2}}}/\sqrt{2\pi\sigma^{2}}$ 0.9991
I.6.20b $f=e^{-\frac{(\theta-\theta_{1})^{2}}{2\sigma^{2}}}/\sqrt{2\pi\sigma^{2}}$ 0.9881
I.8.14 $d=\sqrt{(x_{2}-x_{1})^{2}+(y_{2}-y_{1})^{2}}$ 0.9024
I.9.18 $F=\frac{Gm_{1}m_{2}}{(x_{2}-x_{1})^{2}+(y_{2}-y_{1})^{2}+(z_{2}-z_{1})^{2}}$ 0.9926
I.10.7 $F=\frac{Gm_{1}m_{2}}{(x_{2}-x_{1})^{2}+(y_{2}-y_{1})^{2}+(z_{2}-z_{1})^{2}}$ 0.9872
@@
I.12.11 $F=\mathcal{Q}(E_{f}+Bv\sin\theta)$ 0.9999
I.13.4 $K=\frac{1}{2}m(v^{2}+u^{2}+w^{2})$ 0.9982
I.13.12 $U=Gm_{1}m_{2}(\frac{1}{r_{2}}-\frac{1}{r_{1}})$ 1.0
I.14.3 $U=mgz$ 1.0
I.14.4 $U=\frac{k_{spring}x^{2}}{2}$ 0.9839
I.15.3x $x_{1}=\frac{x-ut}{\sqrt{1-u^{2}/c^{2}}}$ 0.9793
I.15.3t $t_{1}=\frac{t-ux/c^{2}}{\sqrt{1-u^{2}/c^{2}}}$ 0.9638
I.15.10 $p=\frac{m_{0}v}{\sqrt{1-v^{2}/c^{2}}}$ 0.9919
@@
I.18.4 $r=\frac{m_{1}r_{1}+m_{2}r_{2}}{m_{1}+m_{2}}$ 0.9794
I.18.12 $\tau=rF\sin\theta$ 0.9999
I.18.16 $L=mrv\sin\theta$ 0.9999
I.24.6 $E=\frac{1}{4}m(\omega^{2}+\omega_{0}^{2})x^{2}$ 0.9986
@@
I.26.2 $\theta_{1}=\arcsin(n\sin\theta_{2})$ 0.9999
I.27.6 $f_{f}$ $=\frac{1}{\frac{1}{d_{1}}+\frac{n}{d_{2}}}$ 0.9895
I.29.4 $k=\frac{\omega}{c}$ 1.0
I.29.16 $x=\sqrt{x_{1}^{2}+x_{2}^{2}-2x_{1}x_{2}\cos(\theta_{1}-\theta_{2})}$ 0.9828
I.30.3 $I_{*}=I_{*_{0}}\frac{\sin^{2}(n\theta/2)}{\sin^{2}(\theta/2)}$ 0.9912
I.30.5 $\theta=\arcsin(\frac{\lambda}{nd})$ 0.9822
I.32.5 $P=\frac{q^{2}a^{2}}{6\pi\epsilon c^{3}}$ 0.9932
I.32.17 $P=(\frac{1}{2}\epsilon cE_{f}^{2})(8\pi r^{2}/3)(\omega^{4}/(\omega^{2}-\omega_{0}^{2})^{2})$ 0.9817
I.34.8 $\omega=\frac{qvB}{p}$ 1.0
I.34.10 $\omega=\frac{\omega_{0}}{1-v/c}$ 0.9917
I.34.14 $\omega=\frac{1+v/c}{\sqrt{1-v^{2}/c^{2}}}\omega_{0}$ 0.9992
@@
I.39.11 $E=\frac{1}{\gamma-1}p_{F}V$ 0.9914
I.39.22 $P_{F}=\frac{nk_{b}T}{V}$ 0.9976
I.40.1 $n=n_{0}e^{-\frac{mgx}{k_{b}T}}$ 0.9816
I.41.16 $L_{rad}=\frac{\hbar\omega^{3}}{\pi^{2}c^{2}(e^{\frac{\hbar\omega}{k_{b}T}}-1)}$ 0.9213
I.43.16 $v=\frac{\mu_{drift}qV_{e}}{d}$ 0.9981
@@
I.43.43 $\kappa=\frac{1}{\gamma-1}\frac{k_{b}v}{A}$ 0.9215
I.44.4 $E=nk_{b}T\ln(\frac{V_{2}}{V_{1}})$ 0.8017
I.47.23 $c=\sqrt{\frac{\gamma pr}{\rho}}$ 0.9733
I.48.20 $E=\frac{mc^{2}}{\sqrt{1-v^{2}/c^{2}}}$ 0.8629
@@
Table 5: Tested Feynman Equations, part 1.

### v5

@@ 4.4.3 Ablation Study of Modal Alignment
Black-box 0.9937 0.9655 0.9028
Average 0.9934 0.9633 0.9328
Figure 7: Effect of dataset size on fitting performance. From the above plots, It can be found that as the size of the training data increases, the performance of the model on each dataset continues to improve. When the data size reaches 10M, the model has achieved satisfactory results on the selected data sets, and the average $R^{2}$ on each data set is greater than 0.98.
The specific experimental results are shown in Table LABEL:tab:1. Under the same conditions of parameters $\theta$, $\lambda_{1}$ and $\lambda_{2}$, the best overall performance is achieved at $\lambda_{3}$ = 1.0, which is much better than the performance at $\lambda_{3}$ = 0.0. It is not difficult to find that, as the weight $\lambda_{3}$ of contrastive loss ($\mathcal{L_{CON}}$) decreases, the performance ($R^{2}$) of MMSR also decreases gradually. This experimental result fully illustrates that modal alignment positively affects the performance of the MMSR model. This beneficial effect arises from the fact that the contrastive loss can bring the matched data features and the skeleton features closer together in the feature space, and more importantly, it keeps the non-matched data features and the skeleton features farther away from each other. This results in better integration of the latter two features.
In summary, from the above ablation experiments, we can conclude that modality alignment plays a positive role for symbolic regression tasks.
Figure 8: (a) represents joint training. (b) represents separate training. Comparison plots of the two contrastive learning introduction strategies. From the diagram above, we can intuitively see the difference between the Joint training strategy and the Separate training strategy. The Joint training strategy trains the three modules together, while the Separate training strategy trains the two encoders and then the decoder.
Table 3: The results of ablation experiments examining the effect of additional constant encoding.
Dataset $R^{2}$
Hyperparameters $\theta=0.07$ $\theta=0.07$ $\theta=0.07$
$\lambda_{1}=1.0$ $\lambda_{1}=1.0$ $\lambda_{1}=1.0$
$\lambda_{2}=1.0$ $\lambda_{2}=0.1$ $\lambda_{2}=0.0$
$\lambda_{3}=1.0$ $\lambda_{3}=1.0$ $\lambda_{3}=1.0$
Nguyen 0.9999 0.9936 0.9226
Keijzer 0.9983 0.9435 0.8935
Korns 0.9982 0.9437 0.9025
Constant 0.9986 0.9753 0.9382
Livermore 0.9844 0.9346 0.8633
Vladislavleva 0.9862 0.9373 0.8824
R 0.9924 0.9836 0.9146
Jin 0.9943 0.9917 0.9538
Neat 0.9972 0.9626 0.9248
Others 0.9988 0.9935 0.9736
Feynman 0.9913 0.9625 0.9027
Strogatz 0.9819 0.9264 0.8552
Black-box 0.9937 0.9482 0.8636
Average 0.9934 0.9613 0.9069
4.4.4 Ablation Study of Constant Encoding
To assess the impact of incorporating additional constant encoding on the performance of the algorithm, we conducted ablation experiments on the $\mathcal{L_{MSE}}$ component of the loss function $\mathcal{L}$, where $\lambda_{2}$=0 represents the absence of constant encoding. The ablation experiments also consisted of three groups. In each group, the settings of the parameters $\theta$, $\lambda_{1}$ and $\lambda_{3}$ were kept consistent, where $\theta$ = 0.07, $\lambda_{1}$ = $\lambda_{3}$ =1. And the parameter $\lambda_{2}$ was set to 0.0, 0.1 and 1.0 respectively. The experimental results can be found in Table LABEL:tab:lmse. The results indicate that adding constant encoding significantly enhances the performance of the MMSR algorithm. By assigning a unique label to each data set, training becomes more straightforward. For instance, in NeSymReS, the sequence [sin, *, C, x] could represent functions such as sin(4x), sin(4.3x), etc. Despite these functions having a similar structural appearance, they can exhibit considerable differences. Introduction of constant encoding facilitates a one-to-one mapping between data instances and their corresponding expressions, effectively reducing the ambiguity of multiple data sets correlating to the same symbolic sequence, thereby aiding in network training. Moreover, dedicating a unique encoding to constants addresses issues where expressions might share high token-level similarities but differ significantly in achieving specific equation-driven goals like fitting accuracy. This issue typically arises when relying solely on cross-entropy loss between the generated and actual sequences.
4.4.5 Effect of the Size of the Training Data
@@
To test the impact of pre-training data size on model test performance, we train our MMSR model on increasingly larger datasets. More specifically, we use datasets consisting of 10K, 100K, 1M, 5M, and 10M samples to train a model respectively. Except for the amount of data, every other aspect of training is the same as described above, and all are trained for the same epochs. From the trend in Fig. 7, we can see that the performance of the model has been greatly improved as the data set continues to grow. We can see that when the data size reaches 5M, the average $R^{2}$ has reached more than 0.9, and when the data size reaches 10M, The model achieves an average $R^{2}$ of over 0.98 on each dataset. Although the growth rate of average $R^{2}$ is relatively slow with the increase of data, it is still increasing. Therefore, it can be inferred that as we continue to increase the size of the training data, the performance of the model will still reach higher levels.
4.5 Effect of the Introduction Strategy of Contrastive Learning
It is very important to use contrastive learning for modality alignment in multi-modality. During training, there are two strategies for introducing contrastive learning: the joint training strategy and the separate training strategy. A comparison of these two strategies is shown in Fig.8.
  • 1.
@@ 4.5 Effect of the Introduction Strategy of Contrastive Learning
    Joint training strategy: The contrast loss is used as a part of the total loss, and the feature extraction module and the feature fusion module are trained at the same time.
  • 2.
@@
    Separate training strategy: First, train a feature extraction module with contrastive learning, freeze it, and then train a feature fusion module separately.
Figure 9: Performance comparison of joint training and separate training with different training set sizes. From the above plots, it can be seen that the algorithm’s performance increases as the training data size increases. However, under the same data size, the model using the joint training strategy is better than the model using the separate training strategy. And the larger the amount of data, the more obvious this trend.
@@
To prove the superiority of the proposed joint training strategy, we conduct comparative experiments on all datasets. From Fig.9, it can be found that the joint training strategy is better than the separate training strategy in terms of effect. As the size of the dataset increases, this advantage becomes more obvious.
We posit that the underlying reason for this phenomenon lies in the implementation of the joint training strategy, wherein contrastive loss ( $\mathcal{L_{CON}}$ ) is simultaneously optimized alongside cross-entropy loss ($\mathcal{L_{CE}}$) and MSE loss ( $\mathcal{L_{MSE}}$ ). During the training process, the feature extraction module and the feature fusion module are inherently interdependent. This interrelation necessitates that the feature extraction module not only brings the extracted data and skeleton features closer together in the feature space but also ensures that the extracted features are optimally aligned with the requirements of the subsequent feature fusion module. Consequently, this synergy enables a reduction in overall loss. In summary, the joint training strategy facilitates a more effective interaction between each module, enhancing their collaborative performance during concurrent training.
4.6 Model Complexity Evaluation
We measure the complexity of the algorithm in terms of the number of model parameters and the time complexity. In this paper, the transformer of all models uses 16 attention heads and an embedding dimension of 512, and the number of decoder layers is 16. (Note that for MMSR, the first four layers are skeleton encoders. No cross-attention is employed. The last four layers are expression decoders). Each method differs in the data encoder, but there is little difference in the overall number of parameters of the model, and the specific parameter scale is shown in Table 4(Number Size of Parameters). We counted the average inference time for each algorithm on the Nguyen dataset. We test each expression 20 times, record the time it takes to infer it, and then average it overall. The specific results are shown in Table 4(Average Time of inference.
From the comparison of model complexity, it is observable that, aside from the TPSR method having the longest average running time, the proposed MMSR achieves a higher $R^{2}$ compared to other methods under the same conditions of parameter quantity and execution time, demonstrating superior performance.
Table 4: Complexity evaluation of MMSR and the state-of-the-art methods.
Methods Number Size Average Time Average $R^{2}$
of Parameters of inference of test dataset
MMSR 86M 26.6s 0.9999
TPSR[25] 86M 86.3s 0.9948
End2End[56] 86M 24.4s 0.8814
NeSymReS[19] 84M 27.3s 0.8568
SymbolicGPT[18] 88M 28.4s 0.6713
@@ 5 Discussion and Conclusion
In this paper, we propose a symbolic regression algorithm, MMSR. We solve the SR problem as a purely multi-modal problem and obtain decent experimental results. This also proves that it is practical to treat the SR problem as a multi-modal task. Specifically, we treat the data and expression skeleton as two different modalities (e.g. image and text). For the data part, we use SetTransformer for the feature extractor, and for the expression skeleton part, we use the first few layers of the transformer as the feature extractor. Specifically, we introduce contrastive learning for modal feature alignment for better feature fusion. The ablation experiments on more than a dozen datasets show that the introduction of contrastive learning has a great improvement in the performance of the algorithm. In addition, we adopt the One-step scheme(the contrastive loss is trained along with other losses during training) in MMSR. Our experiments show that the One-step scheme is more competitive than the Tow-step scheme, which first trains a feature extraction module with contrastive loss and then trains the feature fusion module separately. In MMSR, alleviating expressions may exhibit high token-level similarities but are suboptimal with respect to equation-specific objectives such as fitting accuracy. Our decoder not only generates the sequence of symbols that make up the skeleton of an expression but also directly generates a set of constants. This set of constants is the same length as the symbol sequence, we take the constant with the same index as the placeholder ’C’ as the value at ’C’, and then the constants are further refined with BFGS to obtain the final expression.
Our algorithm is expected to open new application prospects of multi-modal technology. It promotes more multi-modal correlation techniques to be applied to the field of symbolic regression. We can even use symbolic regression techniques to re-feed multi-modal domains.
@@
Next, we will continue to explore the deep fusion of multi-modal techniques and symbolic regression techniques. It also attempts to solve some of the existing problems of MMSR mentioned above.
@@ Acknowledgements
This work was supported by the National Natural Science Foundation of China under Grant 92370117, and CAS Project for Young Scientists in Basic Research under Grant YSBR-090.
Appendix A Appendix: Detailed Settings of hyperparameters during training and inference
Table A.1: Hyperparameters of SetTransformer
hyperparameters Numerical value
N_p 0
activation ‘relu’
bit16 True
dec_layers 5
dec_pf_dim 512
dim_hidden 512
dim_input 3
dropout 0
input_normalization False
length_eq 60
linear False
ln True
lr 0.0001
mean 0.5
n_l_enc 5
norm True
num_features 20
num_heads 8
num_inds 50
output_dim 60
sinuisodal_embeddings False
src_pad_idx 0
std 0.5
trg_pad_idx 0
Appendix B Appendix: Detailed Settings of the hyperparameters of MMSR
Table B.1: Hyperparameters of MMSR
hyperparameters Numerical value
$\theta$ 0.07
$\lambda_{1}$ 1.0
$\lambda_{2}$ 1.0
$\lambda_{3}$ 1.0
@@ Appendix C Appendix: Test data in detail
Table C.1,C.2,C.3 shows in detail the expression forms of the data set used in the experiment, as well as the sampling range and sampling number. Some specific presentation rules are described below
  • 1.
@@
Table C.1: Specific formula form and value range of the three data sets Nguyen, Korns, and Jin.
@@
Nguyen-2 $4x_{1}^{4}+3x_{1}^{3}+2x_{1}^{2}+x$ U$(-1,1,20)$
Nguyen-5 $\sin(x_{1}^{2})\cos(x)-2$ U$(-1,1,20)$
Nguyen-8 $3]{x}$ U$(0,4,20)$
Nguyen-8′′ $3]{x_{1}^{2}}$ U$(0,4,20)$
@@
Table C.2: Specific formula form and value range of the three data sets neat, Keijzer, and Livermore.
@@
Table C.3: Specific formula form and value range of the three data sets Vladislavleva and others.
@@
The empirical results from our investigation unequivocally affirm that MMSR possesses an exceptional ability to discern the underlying mathematical expressions from a constrained sample size. Notably, the $R^{2}$ values achieved were above 0.99 for a predominant portion of the equations, underscoring the algorithm’s remarkable accuracy in fitting these expressions. These findings decisively position MMSR as a potent tool for addressing complex problems within the domains of physics and mathematics. The broader implications of our study suggest that MMSR holds considerable promise for a wide range of applications across different fields. Detailed experimental results are presented in Table D.1 and Table D.2.
@@
I.6.20a $f=e^{-\theta^{2}/2}/\sqrt{2\pi}$ 0.9998
I.6.20 $f=e^{-\frac{\theta^{2}}{2\sigma^{2}}}/\sqrt{2\pi\sigma^{2}}$ 0.9982
I.6.20b $f=e^{-\frac{(\theta-\theta_{1})^{2}}{2\sigma^{2}}}/\sqrt{2\pi\sigma^{2}}$ 0.9901
I.8.14 $d=\sqrt{(x_{2}-x_{1})^{2}+(y_{2}-y_{1})^{2}}$ 0.9213
I.9.18 $F=\frac{Gm_{1}m_{2}}{(x_{2}-x_{1})^{2}+(y_{2}-y_{1})^{2}+(z_{2}-z_{1})^{2}}$ 0.9928
I.10.7 $F=\frac{Gm_{1}m_{2}}{(x_{2}-x_{1})^{2}+(y_{2}-y_{1})^{2}+(z_{2}-z_{1})^{2}}$ 0.9901
@@
I.12.11 $F=\mathcal{Q}(E_{f}+Bv\sin\theta)$ 0.9989
I.13.4 $K=\frac{1}{2}m(v^{2}+u^{2}+w^{2})$ 0.9991
I.13.12 $U=Gm_{1}m_{2}(\frac{1}{r_{2}}-\frac{1}{r_{1}})$ 0.9993
I.14.3 $U=mgz$ 0.9999
I.14.4 $U=\frac{k_{spring}x^{2}}{2}$ 0.9903
I.15.3x $x_{1}=\frac{x-ut}{\sqrt{1-u^{2}/c^{2}}}$ 0.9802
I.15.3t $t_{1}=\frac{t-ux/c^{2}}{\sqrt{1-u^{2}/c^{2}}}$ 0.9729
I.15.10 $p=\frac{m_{0}v}{\sqrt{1-v^{2}/c^{2}}}$ 0.99983
@@
I.18.4 $r=\frac{m_{1}r_{1}+m_{2}r_{2}}{m_{1}+m_{2}}$ 0.9819
I.18.12 $\tau=rF\sin\theta$ 0.9995
I.18.16 $L=mrv\sin\theta$ 0.9939
I.24.6 $E=\frac{1}{4}m(\omega^{2}+\omega_{0}^{2})x^{2}$ 0.9971
@@
I.26.2 $\theta_{1}=\arcsin(n\sin\theta_{2})$ 0.9983
I.27.6 $f_{f}$ $=\frac{1}{\frac{1}{d_{1}}+\frac{n}{d_{2}}}$ 0.9992
I.29.4 $k=\frac{\omega}{c}$ 0.9998
I.29.16 $x=\sqrt{x_{1}^{2}+x_{2}^{2}-2x_{1}x_{2}\cos(\theta_{1}-\theta_{2})}$ 0.9923
I.30.3 $I_{*}=I_{*_{0}}\frac{\sin^{2}(n\theta/2)}{\sin^{2}(\theta/2)}$ 0.9936
I.30.5 $\theta=\arcsin(\frac{\lambda}{nd})$ 0.9926
I.32.5 $P=\frac{q^{2}a^{2}}{6\pi\epsilon c^{3}}$ 0.9992
I.32.17 $P=(\frac{1}{2}\epsilon cE_{f}^{2})(8\pi r^{2}/3)(\omega^{4}/(\omega^{2}-\omega_{0}^{2})^{2})$ 0.9849
I.34.8 $\omega=\frac{qvB}{p}$ 0.9999
I.34.10 $\omega=\frac{\omega_{0}}{1-v/c}$ 0.9915
I.34.14 $\omega=\frac{1+v/c}{\sqrt{1-v^{2}/c^{2}}}\omega_{0}$ 0.9982
@@
I.39.11 $E=\frac{1}{\gamma-1}p_{F}V$ 0.9993
I.39.22 $P_{F}=\frac{nk_{b}T}{V}$ 0.9981
I.40.1 $n=n_{0}e^{-\frac{mgx}{k_{b}T}}$ 0.9868
I.41.16 $L_{rad}=\frac{\hbar\omega^{3}}{\pi^{2}c^{2}(e^{\frac{\hbar\omega}{k_{b}T}}-1)}$ 0.9453
I.43.16 $v=\frac{\mu_{drift}qV_{e}}{d}$ 0.9941
@@
I.43.43 $\kappa=\frac{1}{\gamma-1}\frac{k_{b}v}{A}$ 0.9513
I.44.4 $E=nk_{b}T\ln(\frac{V_{2}}{V_{1}})$ 0.8616
I.47.23 $c=\sqrt{\frac{\gamma pr}{\rho}}$ 0.9639
I.48.20 $E=\frac{mc^{2}}{\sqrt{1-v^{2}/c^{2}}}$ 0.8935
@@
Table D.1: Tested Feynman Equations, part 1.